OpenStudy (anonymous):

clear fractions or decimals , solve and check 3.7x + 45.5 = 4.7 - 6.5x

OpenStudy (anonymous):

What do you mean when you say "clear fractions or decimals"

OpenStudy (anonymous):

3.7x + 45.5 = 4.7 - 6.5x +6.5x +6.5x 10.2x+45.5=4.7 -45.5 -45.5 10.2x=-40.8 divide all by 10.2 and you get x=-4

OpenStudy (anonymous):

If you want to check your work then plug your "Y" back in like 3.7(-4)+ 45.5 = 4.7 - 6.5(-4) 30.7=30.7 so you are right! @ChmE I'm pretty sure that just means make sure that you do not have any fractions or decimals left. That's why it says "clear".